If an amount of money in P in pesos is invested at " r " percent compounded annually, it will grow to an amount A = P(1 + r)^2 in two years. Suppose Miss Madrigal wants her money amounting to Php 200,000 to grow Php 228, 980 in two years. At what rate must she invest her money
Answer by Theo(13342)

a = p(1+r)^2
a = 228980
p = 200000
replace a and p in the equation to get:
228980 = 200000(1+r)^2
divide both sides by 200000 to get:
228980/200000 = (1+r)^2
simplify to get:
1.14449 = (1+r)^2
take the square root of both sides to get:
1.07 = 1+r
subtract 1 from both sides to get:
r = .07
the rate she needs to invest her money at is 7%.